The proton And the eletron are 212 pm apart in the plane of the paper. Find their eletric dipole moment
Answer #1

q= charge = 1.6*10-19 C

d= distance between the charges = 212*10-12 m

dipole moment = q*d

= 1.6*10-19 * 212*10-12

   = 3.392*10-29 C-m